[Jack is 36 years old. He went to a birthday party for someone in his family named Alicia. When he was there, he realized that his age is AN INTEGER-MULTIPLE of Alicia's age. Find all the possible ages that Alicia could be.]
Answer by TimothyLamb(4379)

---
possible ages of alicia
36/1 = 36
36/2 = 18
36/3 = 12
36/4 = 9
36/6 = 6
36/9 = 4
36/12 = 3
36/36 = 1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php